# DocFusion: A Unified Framework for Document Parsing Tasks

### v1

@@ Abstract
Document parsing is essential for analyzing complex document structures and extracting fine-grained information, supporting numerous downstream applications. However, existing methods often require integrating multiple independent models to handle various parsing tasks, leading to high complexity and maintenance overhead. To address this, we propose DocFusion, a lightweight generative model with only 0.28B parameters. It unifies task representations and achieves collaborative training through an improved objective function. Experiments reveal and leverage the mutually beneficial interaction among recognition tasks, and integrating recognition data significantly enhances detection performance. The final results demonstrate that DocFusion achieves state-of-the-art (SOTA) performance across four key tasks.
@@
Mingxu Chai1,2, Ziyu Shen1, Chong Zhang1, Yue Zhang1, Xiao Wang1, Shihan Dou1, Jihua Kang1, Jiazheng Zhang 1, Qi Zhang1 1Fudan University, 2 Shanghai Innovation Institute
@@ 1 Introduction
Document parsing plays a crucial role in extracting structured data from complex documents, serving as a foundational technology for downstream applications. It is particularly important in Retrieval-Augmented Generation (RAG) workflows Ren et al. (2023); Zhang et al. (2022), where extracting organized and contextually rich information from documents can significantly enhance the performance of large language models (LLMs) Jiang et al. (2023); Zhao et al. (2024a); Gao et al. (2024). However, information in real-world documents is often embedded in complex structures, such as hierarchical layouts, mathematical expressions, and tables, which makes automatic parsing substantially challenging.
Figure 1: Pipeline systems integrate multiple modules into a Framework. In contrast, DocFusion incorporates multiple functionalities within a single model.
@@
To address these issues, research on document parsing has primarily focused on four key tasks: document layout analysis (DLA), mathematical expression recognition (MER), table recognition (TR), and optical character recognition (OCR). Existing methods can be categorized into two main approaches: multi-module pipeline systems and end-to-end page-level OCR models. As shown in Figure 1, multi-module pipeline systems decompose document parsing tasks into independent modules, allowing each module to adopt the most suitable model. For example, DocLayout-YOLO Zhao et al. (2024c) has demonstrated excellent performance in DLA, while UniMERNet Wang et al. (2024a) achieves SOTA results in MER. Although this approach improves performance for specific tasks, integrating multiple models into a single system increases overall complexity. Moreover, these systems fail to fully exploit task-level collaboration, leading to inefficiencies in parameter usage. In contrast, end-to-end page-level OCR models, such as Nougat Blecher et al. (2023) and GOT Wei et al. (2024), can seamlessly integrate multiple recognition tasks. While the outputs of these models demonstrate a well-organized logical structure, the models lack the ability of DLA to generate bounding boxes (bboxes) for layout elements. As a result, they cannot preserve the spatial relationships between documents and their corresponding layouts. This limitation is critical in RAG workflows, where preserving spatial relationships is essential for achieving interpretability. The absence of DLA also affects single-task applications such as MER and TR, which depend on accurate layout analysis for reliable results. These limitations highlight the urgent need for an approach to reduce system complexity and integrate multiple tasks. Therefore, we aim to develop a model capable of simultaneously handling DLA, MER, OCR, and TR tasks.
In this paper, we propose DocFusion, a unified generative multi-task model designed to address four key document parsing tasks. DocFusion leverages multi-task collaboration to achieve comprehensive optimization in document parsing. To handle complex layouts, we introduce Dual Attention Ding et al. (2022), which combines spatial and channel information interactions. This mechanism enhances DocFusion’s ability to process complex tasks with greater accuracy. To address the challenge of loss convergence in detection tasks (DLA) within a generative framework, we design a specialized objective function. The challenge arises from the conflict between the continuous nature of coordinate data and the discrete nature of token generation. Our objective function applies a one-dimensional convolution to smooth the discrete generation distribution. This approach significantly accelerates loss convergence and enables efficient joint training.
Experimental results demonstrate that DocFusion achieves leading performance across all four tasks. Additionally, the recognition tasks mutually enhance each other’s performance, leading to overall improvements compared to single-task setups. Notably, OCR improves DLA by providing enriched textual context, enabling more precise layout analysis. Further experiments validate the effectiveness of the improved objective function, demonstrating its key role in enabling task collaboration and performance gains.
@@
  • We propose DocFusion, a unified generative multi-task model that standardizes task formulations and achieves SOTA performance across four key document parsing tasks: DLA, MER, TR, and OCR.
  • Experimental results demonstrate that incorporating multi-task data significantly outperforms single-task setups, providing insights into the benefits of multi-task learning.
  • We propose an improved objective function to directly address the conflict between the continuous nature of coordinate data and the discrete nature of token generation in detection tasks within the generative framework.
  • We constructed a large-scale dataset containing 1.5M LaTeX-annotated math expressions and 100K tables, standardized for consistency, providing a valuable resource for advancing document parsing research.
@@
Tool Size Type DLA MER TR OCR
UniMER (2024a) 325M M
DocLayout(2024c) 20M M
StructTable (2024) 938M M
ViTLP (2024) 253M M
Nougat (2023) 350M M
GOT (2024) 580M M
open-parse (2024) - S
LlamaParse (2024) - S
DeepDoc (2024) - S
MinerU (2024) - S
DocFusion 289M M
Table 1: Capabilities of document parsing tools. Type: M represents a model, while S denotes a system. DLA: Document Layout Analysis. MER: Math Expression Recognition. TR: Table Recognition. OCR: Optical Character Recognition.
@@ 2 Related Work
Document Parsing Models. Document parsing models have seen remarkable progress across various tasks. DLA has evolved from vision-based methods Wick and Puppe (2018); Bao et al. (2021) to multimodal approaches integrating textual features Xu et al. (2022); Huang et al. (2022). OCR has transitioned from template matching Smith (2007) to deep learning-based solutions Bušta et al. (2017); Chen et al. (2021); Mosbah et al. (2024). MER progressed from symbol segmentation Miller and Viola (1998) to CNN-RNN hybrids Le et al. (2019) and Transformer-based models Wang et al. (2024a). Similarly, TR now employs methods like grid segmentation and image-to-sequence techniques to reconstruct structured data Qasim et al. (2019); Huang et al. (2023); Xia et al. (2024). Page-level end-to-end OCR models like Nougat Blecher et al. (2023) and GOT Wei et al. (2024) simplify workflows by integrating multi recognition tasks.
Modular Pipeline Systems. The advancements in task-specific models have driven the development of modular pipeline systems, which process complex document structures through specialized modules. For instance, Open-ParseFilimonov (2024) performs well in incrementally parsing complex layouts but lacks support for MER. Other systems, such as DeepDocYu (2024) and Llama-ParseLiu (2024), extend the scope of modular pipelines to handle more diverse tasks. In particular, MinerUZhao et al. (2024b) stands out by supporting advanced features such as complex layout parsing and Markdown conversion. However, despite their flexibility, modular systems face significant challenges in practical deployment. The variability in environmental dependencies between modules increases the complexity of maintenance. Furthermore, tasks that could be efficiently handled by a single module are often divided among multiple modules, leading to unnecessary system overhead. These limitations highlight the need for more unified and efficient frameworks to address the growing demands of document parsing.
3 DocFusion
We introduce the model architecture (3.1) and explain how detection tasks are integrated into the generative framework. Then, we discuss the challenges (3.2) of detection tasks within this framework. Next, we explain the improved objective function (3.3)
@@ 3.1 Architecture
As shown in Figure 2, the architecture of DocFusion consists of three main components: a vision encoder, a text embedding layer, and a Transformer decoder. Since the task instructions are limited and predefined, no Transformer encoder is included; instead, task-specific prompts are directly embedded, simplifying the architecture. To unify the representation of object detection and text recognition tasks, we adopt a coordinate quantization representation Xiao et al. (2023). Specifically, images are quantized into a fixed resolution (e.g., 1000×1000), and coordinates are represented as discrete tokens (e.g., <loc_1>, <loc_2>, …, <loc_1000>). This approach enables the use of a unified regression framework for detection tasks, simplifying multi-task integration. To address the challenges posed by densely structured content, the vision encoder incorporates a Dual Attention mechanism Ding et al. (2022), which captures interactions across channel and spatial dimensions, enhancing feature extraction for intricate document layouts. Additionally, the traditional feed-forward network (FFN) is removed, reducing both parameter count and computational cost, further improving model efficiency. The vision encoder processes input images $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ into visual features, flattened as token embeddings $\mathbf{V}\in\mathbb{R}^{N_{v}\times D_{v}}$. These embeddings are transformed for compatibility with task-specific prompt embeddings $\mathbf{T}{\text{prompt}}\in\mathbb{R}^{N_{t}\times D}$. The combined input $\mathbf{X}=[\mathbf{V^{\prime}};\mathbf{T}{\text{prompt}}]$ is then passed to the Transformer decoder to generate predictions. By integrating Dual Attention, coordinate quantization, and optimizing its architecture, DocFusion efficiently handles complex document parsing tasks with high precision and computational efficiency.
Figure 3: The Softmax distribution of logits for a target token and its neighboring tokens after the loss has stabilized.
@@ 3.2 Challenges and Motivations
While representing object detection as text regression enables joint training of layout analysis and page element recognition under a unified cross-entropy-based framework, it inherently forces continuous coordinates into discrete token spaces. This mismatch creates several challenges, especially in fine-tuning small coordinate adjustments, where the model struggles to produce accurate gradients, reducing training stability. As shown in Figure 3, small unavoidable deviations in coordinate labels smooth out the softmax distribution, preventing the target token’s probability from forming a sharp peak. This makes it harder for the model to escape local optima and limits its learning capacity. Additionally, traditional cross-entropy loss, which is designed for discrete classification tasks, does not handle continuous changes well, further increasing inaccuracies during training.
@@
3.3 Objective function
To address these challenges, we propose an improved objective function that applies a one-dimensional convolution over the probability distribution, refining the model’s sensitivity to small coordinate changes while preserving the discrete treatment of cross-entropy. This approach helps alleviate the mismatch between discrete tokens and continuous coordinates, improves gradient quality, and prevents the coordinate prediction task from dominating the optimization process. In doing so, it enhances localization accuracy, supports stable multi-task training, and achieves better alignment with the desired properties identified in the motivating considerations.
@@ 3.3 Objective function
Next, a one-dimensional convolution kernel $\mathbf{K}\in\mathbb{R}^{1\times 1\times k}$ is constructed based on a Gaussian distribution, where $k$ is the kernel size (an odd integer greater than 1), and $\sigma$ is the standard deviation. The kernel weights are computed as:
$\mathbf{K}_{i}=\exp\left(-\frac{(i-\frac{k-1}{2})^{2}}{2\sigma^{2}}\right)$ (3)
The kernel is then applied to $\mathbf{P}^{\prime}$ via one-dimensional convolution along the vocabulary dimension:
$\mathbf{C}=\text{conv1d}(\mathbf{P}^{\prime},\mathbf{K},\text{padding}=\frac{k-1}{2})$ (4)
This convolution preserves the size of the input and output tensors.The convolution result $\mathbf{C}$ is integrated back into the original probability distribution $\mathbf{P}$ within the index range $[s,e]$, while retaining the original probabilities outside this range:
@@
4 Dataset Construction and Refinement
In this section, we briefly describe the reconstruction of the DLA dataset and the collection of the MER, TR, and OCR datasets, with more detailed information provided in the appendix.
DLA Dataset. DocLaynet Pfitzmann et al. (2022) was chosen for its comprehensive layout annotations, but its formula annotations, where content and numbering share the same bounding box, introduce noise for MER tasks. To address this, we re-extracted formulas from arXiv LaTeX files, trained a lightweight model to re-annotate the pages with manual verification.
MER Dataset. The UniMER-1M Wang et al. (2024a) has significantly advanced MER research but contains many redundant spaces in LaTeX code. Although some spaces are syntactically necessary, most are unnecessary, increasing output length and computational overhead. To address this, we constructed a new dataset by extracting content from LaTeX files, normalizing style variations and verifying accuracy through re-rendering. Models trained on our dataset produce LaTeX code that is approximately 34.2% shorter for complex expressions and 37.5% shorter for simple expressions on the UniMER-1M test set, demonstrating improved efficiency and performance.
OCR Dataset. The dataset also sourced from DocLaynet Pfitzmann et al. (2022), provides detailed layout and character annotations. We extracted cropped images for each layout element and paired them with corresponding character-level text annotations.

### v2

@@ Abstract
Document parsing involves layout element detection and recognition, essential for extracting information. However, existing methods often employ multiple models for these tasks, leading to increased system complexity and maintenance overhead. While some models attempt to unify detection and recognition, they often fail to address the intrinsic differences in data representations, thereby limiting performance in document processing. Our research reveals that recognition relies on discrete tokens, whereas detection relies on continuous coordinates, leading to challenges in gradient updates and optimization. To bridge this gap, we propose the Gaussian-Kernel Cross-Entropy Loss (GK-CEL), enabling generative frameworks to handle both tasks simultaneously. Building upon GK-CEL, we propose DocFusion, a unified document parsing model with only 0.28B parameters. Additionally, we construct the DocLatex-1.6M dataset to provide high-quality training support. Experimental results show that DocFusion, equipped with GK-CEL, performs competitively across four core document parsing tasks, validating the effectiveness of our unified approach. The model and datasets are publicly available at: https://github.com/sc22mc/DocFusion
@@
Mingxu Chai1,2thanks: Equal contribution, Ziyu Shen1footnotemark: , Chong Zhang1, Yue Zhang1, Xiao Wang1, Shihan Dou1, Jihua Kang1, Jiazheng Zhang1, Qi Zhang1thanks: Corresponding author 1Computation and Artificial Intelligence Innovative College, Fudan University, Shanghai, China 2Shanghai Innovation Institute, Shanghai, China {qz}@fudan.edu.cn
@@ 1 Introduction
Document parsing plays a significant role in extracting structured data from documents, making it foundational for various downstream applications. For example, in Retrieval-Augmented Generation (RAG) workflows Ren et al. (2023); Zhang et al. (2022), extracting well-organized and contextually rich information from documents can improve the performance of large language models (LLMs) Zhao et al. (2024a); Gao et al. (2024). However, real-world documents often embed information in complex structures, such as hierarchical layouts, mathematical expressions, and tables, which pose considerable challenges for automated parsing.
Tool Type Size DLA MER TR OCR
System
open-parse (2024) -
LlamaParse (2024) -
DeepDoc (2024) -
MinerU (2024) -
Model
DocYOLO(2024c) 20M
ViTLP (2024) 253M
UniMER (2024b) 325M
Nougat (2023) 350M
GOT (2024) 580M
StructTable (2024) 938M
DocFusion(Ours) 289M
Table 1: Capabilities of document parsing tools. Model refers to a single model, while System integrates multiple models. DLA: Document Layout Analysis. MER: Math Expression Recognition. TR: Table Recognition. OCR: Optical Character Recognition. Compare with multi-model systems, DocFusion achieves all four tasks within a single model, requiring only 289M parameters.
@@
Existing methods can be categorized into two main approaches: multi-module pipeline systems and end-to-end page-level OCR models. Multi-module pipeline systems decompose document parsing tasks into independent modules, allowing each module to adopt the best model. For example, DocLayout-YOLO Zhao et al. (2024c) has demonstrated excellent performance in Layout analysis, while UniMERNet Wang et al. (2024b) achieves SOTA results in Math Expression Recognition. Although this approach improves performance for specific tasks, integrating multiple models into a single system increases overall complexity. Moreover, these systems fail to fully exploit task-level collaboration, leading to inefficiencies in parameter usage. In contrast, end-to-end page-level OCR models, such as Nougat Blecher et al. (2023) and GOT Wei et al. (2024), can seamlessly integrate multiple recognition tasks. While the outputs of these models demonstrate a well-organized logical structure, the models lack the ability to generate bounding boxes for layout elements. As a result, they fail to preserve the spatial relationships between documents and their layouts, which is crucial for interpretability in RAG workflows. Additionally, while these models perform well on page-level images, it struggles with specific layout elements, limiting their flexibility in application. To address these issues, this research focused on four key tasks: document layout analysis (DLA), mathematical expression recognition (MER), table recognition (TR), and optical character recognition (OCR).
Several studies have attempted to apply generative frameworks to integrate object detection and content recognition, achieving promising results on natural images Xiao et al. (2023). However, extending such frameworks to document images presents significant challenges due to the inherent structural and representational differences between these domains. Through experiments, we identify the primary issue as the fundamental conflict between the continuous nature of coordinate data and the discrete nature of token generation, which disrupts gradient updates during multi-task training (discussed in Section 3.2). In natural scene images, small deviations in coordinates and text are generally tolerable. However, in document parsing, even minor errors in LaTeX code can critically impact compilation success rates. This imposes stricter accuracy requirements when applying such frameworks to document understanding tasks. To address these challenges, we propose Gaussian-Kernel Cross-Entropy Loss (GK-CEL), an improved objective function designed to mitigate the inconsistencies between discrete and continuous data representations, enhancing the performance of generative frameworks in document parsing.
MER and TR are essential for LaTeX-based document processing, but existing datasets suffer from inconsistent formatting and redundant characters, where different writing styles generate identical compiled outputs, introducing noise that hinders model training (details are provided in Appendix A.3). To address this, we propose DocLatex-1.6M, a large-scale, high-quality dataset that enhances annotation consistency and improves model training efficiency. Experiments demonstrate that DocFusion trained on this dataset outperforms task-specific models with fewer parameters.
@@
  • We propose DocFusion, a unified generative multi-task model that standardizes task formulations and achieves SOTA performance across four key document parsing tasks.
  • We propose GK-CEL to resolve the conflict between continuous coordinate and discrete token in the generative framework, enhancing document parsing and offering a reference for similar frameworks in other domains.
  • Experimental results demonstrate that incorporating multi-task data significantly outperforms single-task setups, providing insights into the benefits of multi-task learning in document parsing.
  • We constructed DocLatex-1.6M, a large-scale, high-quality dataset with 1.5M LaTeX-annotated math expressions and 100K tables, offering a valuable resource for advancing document parsing research.
@@ 2 Related Work
Document Parsing Models. Document parsing models have seen remarkable progress across various tasks. DLA has evolved from vision-based methods Wick and Puppe (2018); Bao et al. (2021) to multimodal approaches integrating textual features Xu et al. (2022); Huang et al. (2022). OCR has transitioned from template matching Smith (2007) to deep learning-based solutions Bušta et al. (2017); Chen et al. (2021); Mosbah et al. (2024). MER progressed from symbol segmentation Miller and Viola (1998) to CNN-RNN hybrids Le et al. (2019) and Transformer-based models Wang et al. (2024b). Similarly, TR now employs methods like grid segmentation and image-to-sequence techniques to reconstruct structured data Qasim et al. (2019); Huang et al. (2023); Xia et al. (2024). Page-level end-to-end OCR models like Nougat Blecher et al. (2023) and GOT Wei et al. (2024) simplify workflows by integrating multi recognition tasks.
Modular Pipeline Systems. The advancements in task-specific models have driven the development of modular pipeline systems, which process complex document structures through specialized modules. For instance, Open-ParseFilimonov (2024) performs well in incrementally parsing complex layouts but lacks support for MER. Other systems, such as DeepDocYu (2024) and Llama-ParseLiu (2024), extend the scope of modular pipelines to handle more diverse tasks. In particular, MinerUZhao et al. (2024b) stands out by supporting advanced features such as complex layout parsing and Markdown conversion.
3 Method
We introduce the model architecture (3.1) and explain how detection tasks are represented into the generative framework. Then, we discuss the challenges (3.2) of detection tasks within this framework. Next, we explain the Gaussian-Kernel Cross-Entropy Loss(3.3)
@@ 3.1 Architecture
As shown in Figure 1, the architecture of DocFusion consists of three main components: a vision encoder, a text embedding layer, and a Transformer decoder. Since the task instructions are limited and predefined, no Transformer encoder is included, task-specific prompts are directly embedded, simplifying the architecture.
To unify the representation of object detection and text recognition tasks, we adopt a coordinate quantization representation Xiao et al. (2023). Specifically, images are quantized into a fixed resolution (e.g., 1000×1000), and coordinates are represented as discrete tokens (e.g., <loc_1>, <loc_2>, …, <loc_1000>). This approach enables the use of a unified generative framework for detection tasks. To address the challenges posed by densely structured content, the vision encoder incorporates a Dual Attention mechanism Ding et al. (2022), which captures interactions across channel and spatial dimensions, enhancing feature extraction for intricate document layouts. Additionally, the traditional feed-forward network (FFN) is removed, reducing both parameter count and computational cost, further improving model efficiency.
The vision encoder processes input images $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ into visual features, flattened as token embeddings $\mathbf{V}\in\mathbb{R}^{N_{v}\times D_{v}}$. These embeddings are projected to $D_{t}$, resulting in $\mathbf{V^{\prime}}\in\mathbb{R}^{N_{v}\times D_{t}}$, to match the task-specific prompt embeddings $\mathbf{T}{\text{prompt}}\in\mathbb{R}^{N_{t}\times D_{t}}$. The combined input $\mathbf{X}=[\mathbf{V^{\prime}};\mathbf{T}{\text{prompt}}]$ is then passed to the Transformer decoder to generate predictions.
Figure 2: The distribution of logits for a target token after the loss has stabilized when using the Common CE Loss.
@@ 3.2 Challenges and Motivations
While representing object detection as text generation enables joint training of layout analysis and page element recognition under a unified cross-entropy-based framework, it inherently forces continuous coordinates into discrete token spaces. This mismatch creates several challenges, especially in fine-tuning small coordinate adjustments, where the model struggles to produce accurate gradients, reducing training stability. As shown in Figure 2, small unavoidable deviations in coordinate labels smooth out the softmax distribution, preventing the target token’s probability from forming a sharp peak. This makes it harder for the model to escape local optima and limits its learning capacity. Additionally, traditional cross-entropy loss, which is designed for discrete classification tasks, does not handle continuous changes well, further increasing inaccuracies during training.
@@
3.3 Gaussian-Kernel Design
To address these challenges, we propose the Gaussian-Kernel Cross-Entropy Loss (GK-CEL). As shown in Figure 3, it applies a one-dimensional convolution with Gaussian-distributed weights over the probability distribution, fine-tuning the model’s sensitivity to small coordinate changes while preserving the discrete treatment of cross-entropy. This approach alleviates the mismatch between discrete tokens and continuous coordinates, improves gradient quality, and prevents the coordinate prediction task from dominating the optimization process. As a result, it enhances localization accuracy and supports stable multi-task training.
Figure 3: Illustration of Gaussian-Kernel Cross-Entropy Loss.
@@ 3.3 Gaussian-Kernel Design
where $i$ represents the batch index, $j$ represents the sequence position, and $k$ represents the vocabulary index.
Next, a one-dimensional convolution kernel $\mathbf{K}\in\mathbb{R}^{1\times 1\times n}$ is constructed based on a Gaussian distribution, where $n$ is the kernel size (an odd integer greater than 1), $\sigma$ is the standard deviation and $p$ represents the position of each element in the convolution kernel, measured as the offset from the center, where the center is located at $\frac{n+1}{2}$. The range of $p\in[1,n]$. The kernel weights of each index are computed as:
$\mathbf{K}_{p}=\exp\left(-\frac{(p-\frac{n+1}{2})^{2}}{2\sigma^{2}}\right)$ (3)
The kernel is then applied to $\mathbf{P}^{\prime}$ via one-dimensional convolution:
$\mathbf{C}=\text{conv1d}(\mathbf{P}^{\prime},\mathbf{K})$ (4)
The convolution result $\mathbf{C}$ is integrated back into the original probability distribution $\mathbf{P}$ within the index range $[s,e]$, while retaining the original probabilities outside this range:
@@
where $\mathbf{M}_{ij}$ is a mask matrix that indicates whether the target label at position $(i,j)$ should contribute to the loss calculation.The normalization factor $\mathbf{N}$ is defined as the total number of valid targets.
@@ 4 Experiments
4.1 Training Datasets
In the training phase, the DLA task uses the DocLayNet Pfitzmann et al. (2022) dataset, which contains 80,863 pages from 7 document types and is manually annotated with 11 categories. The images are split into 69,103/6,481/4,999 for training/validation/testing, respectively. The OCR dataset is also sourced from DocLayNet, which offers comprehensive annotations for layout elements and their corresponding text, and is widely regarded as a reliable resource in the academic community. For the TR and MER tasks, we used the DocLatex-1.6M dataset, which was constructed in this work. Additionally, although this work primarily focuses on document images, we introduced the HME100KYuan et al. (2022), a handwritten math expression dataset to enhance the generalization ability of the MER task.
Model size OCR MER TR
[…]
UReader (2023) 7B 38.6 47.3 - - - - -
LLaVA-NeXT (2024) 34B 69.1 27.2 - - - - -
Nougat (2023) 250M 71.6 21.4 - - - - -
TextMonkey (2024) 7B 73.3 21.9 - - - - -
Qwen-VL-MAX (2023) >72B 94.7 3.9 - - - - -
Qwen-VL-OCR (2023) - 95.9 4.1 - - - - -
Pix2tex (2022) - - - 76.5 41.7 95.9 - -
Texify (2023) 312M - - 88.6 71.7 97.8 - -
Mathpix - - - 88.9 79.1 98.3 - -
UniMERNet (2024b) 325M - - 99.0 89.5 99.7 - -
MixTex (2024) 85M - - - - - 46.2 27.4
StructEqTable (2024) 938M - - - - - 90.6 93.2
GOT (2024) 580M 96.8 2.2 87.7 67.3 97.8 86.9 81.6
DocFusion(Ours) 289M 97.4 1.8 98.7 94.2 99.8 92.1 92.5
Table 2: Comparison of DocFusion with other models on three recognition tasks in the document scene. Specifically, DocLaynetPfitzmann et al. (2022) was used for OCR, DocGenomeXia et al. (2024) for TR, and UniMER-1MWang et al. (2024b) for MER. More details on the TR experiments can be found in Appendix B.2. Note: Nougat is primarily designed for full-page recognition and tends to underperform on isolated tables or mathematical expressions.
[…]
YOLOv10m (2024a) 16M 90.1 86.9 88.4 94.3 94.5 94.4 93.6
YOLOv11m (2024) 20M 90.5 87.4 88.9 95.1 94.9 95.0 100.8
YOLO-DocLayout (2024c) 20M 90.9 88.2 89.5 95.5 94.4 95.0 55.2 -
DETR (2020) 41M 84.7 87.1 85.8 92.2 92.0 92.1 17.6
DETR-Deformable (2020) 41M 91.6 87.1 89.3 96.2 95.9 96.0 18.8
DocFusion(Ours) 289M 88.9 87.8 88.4 96.8 96.2 96.4 11.4
Table 3: The performance of the models on DLA, where DocLayNet-Scientific refers to the scientific document subset of DocLayNet. NMS indicates that Non-Maximum Suppression is not required, while Conf means no confidence adjustment is needed. Note: The results of DETR and YOLO-series models in this table are computed at multiple confidence levels, with the highest F1 score selected as the final result. YOLO-DocLayout builds on YOLOv10, which is NMS-free by design. However, due to structural changes, it is unclear whether it can still be fully considered NMS-free.
4.2 Evaluation Metrics
4.2.1 Evaluation Metrics for Recognition
We employ BLEU Papineni et al. (2001) and Edit Distance Levenshtein (1966) to evaluate sequences. Additionally, CDM Wang et al. (2024c) and CSR were used to better assess the quality of LaTeX-based outputs.
BLEU is used for evaluating generated text, measuring n-gram overlap with reference texts.
Edit distance measures the minimum number of operations insertions, deletions, or substitutions required to transform one string into another.
CSR refers to the percentage of generated LaTeX outputs that can be successfully compiled into PDF.
ExpRate Li et al. (2022) measures the proportion of samples where the predicted text matches the reference text without any errors.
CDM evaluates MER by comparing image-rendered expression at the character level with spatial localization, ensuring fairness and accuracy over text-based metrics like BLEU.
4.2.2 Evaluation Metrics for Detection
Since DocFusion adopts a novel approach in the DLA task without relying on confidence scores, we did not use the widely adopted Average Precision but instead focus on the following metrics:
Precision measures the proportion of correctly identified positive instances among all predicted positives.
Recall measures the proportion of correctly identified positive instances among all actual positives.
F1-score balances precision and recall, serving as their harmonic mean.
FPS measures the number of frames processed by the model per second.
Train Dataset OCR MER TR DLA
[…]
Task-Specific 96.7 2.2 98.5 99.8 91.2 92.7 87.8
OCR+DLA 96.1 2.4 - - - - 88.9
OCR+MER+TR 97.1 1.8 98.9 99.9 92.3 94.6 -
4.3 Selection of Baseline Models
For the MER task, we selected UniMERNet Wang et al. (2024b), the current state-of-the-art (SOTA) model, and Texify Paruchuri (2023), which has shown strong competitive performance in recent evaluations. In the OCR task, we compared several models, including the large-scale model TextMonkey Liu et al. (2024) and smaller models such as Nougat Blecher et al. (2023), for a multi-scale evaluation. For the TR task, we evaluated our approach against StructEqTable Xia et al. (2024), one of the most representative models in current Table-to-Sequence methods. In the DLA task, we compared our method with two major object detection frameworks, YOLO and DETR (e.g., DocLayout-YOLO Zhao et al. (2024c), Deformable-DETR Zhu et al. (2020)). Although GOT Wei et al. (2024) is not capable of handling the DLA task, it performs well in the other three recognition tasks, making it a relevant model for comparison.